OpenStudy (anonymous):

Solve for n: 6(n – 2) + 2n = 7n + 3(n – 2) I get positive 3. But the answers are -3,-9, all real solutions, no solution. what am i doing wrong?

OpenStudy (anonymous):

6(n – 2) + 2n = 7n + 3(n – 2) 6n-12+2n=7n+3n-6 8n-12=10n-6 8n-10n=-6+12 -2n=6 multiply both sides by (-1) you will get.. 2n=-6 so n=-3 :)